# Improved real-space parallelizable matrix-product state compression and its application to unitary quantum dynamics simulation

### v1

@@ Appendix B Details of the quantum circuits adopted in the benchmarking simulations
B.2 One-dimensional random quantum circuit (RQC-1D)
@@
where $\sigma^{x}_{r}$, $\sigma^{y}_{r}$, and $\sigma^{z}_{r}$ are Pauli matrices, and $\alpha_{r}^{(l)}$, $\theta_{r}^{(l)}$, and $\phi_{r}^{(l)}$ are uniformly distributed real random numbers. After applying $U_{r}^{(l)}$ to all qubits in the $l$th circuit layer, the quantum entanglement is generated by subsequently applying a layer of controlled-Z (CZ) gates, completing one physical layer of the circuit construction. The CZ gates apply to $(0,1),(2,3),\cdots,(N-3,N-2)$ qubit pairs for odd physical layers and on $(1,2),(3,4),\cdots,(N-2,N-1)$ qubit pairs for even physical layers. The total number $N$ of qubits is restricted to odd, and the total number $D$ of physical circuit layers is restricted to even.
@@
B.3 One-dimensional parametrized quantum circuit (PQC-1D)
@@
B.4 Two-dimensional random quantum circuit (RQC-2D)
RQC-2D is the direct extension of RQC-1D on a 2D square lattice with the number of qubits $N=L_{x}\times L_{y}$. The physical circuit layers are applied in an ABCDABCD $\cdots$ pattern, as illustrated in Fig. 9, where each rectangle indicates the qubits on which a CZ gate acts. Additionally, single-qubit random gates $\{U_{r}^{(l)}(\theta_{ij}^{(l)})\}$ are applied to all qubits in each physical circuit layer, similar to the case of RQC-1D. The number of physical circuit layers is restricted to a multiple of 4. However, there are no specific restrictions on $L_{x}$ and $L_{y}$. In cases when $L_{x}$ ($L_{y}$) is odd, no two-qubit gates are applied on qubits at the left and right (top and bottom) edges in physical circuit layers C and D (A and B), respectively, as shown in Fig. 9 for $L_{x}=L_{y}=5$.
@@ B.4 Two-dimensional random quantum circuit (RQC-2D)
In contrast to 1D quantum circuits, where qubits naturally align in a 1D path suitable for the MPS representation, a mapping of qubits for 2D quantum circuits to a 1D path suitable for the MPS representation must be determined. Here, we choose the path shown in Fig. 10(a) to ensure that two-qubit gates in physical circuit layers A and B act on neighboring qubits in the sense of the MPS 1D path. Under this mapping, the two-qubit gates in physical circuit layers C and D become long-distance gates, necessitating a recompilation of these two physical circuit layers for MPS-based simulations to ensure that gates apply only to neighboring qubits on the MPS 1D path. As an example, the recompiling procedure for the $L_{x}=L_{y}=4$ case is illustrated in Fig. 10(b). Note that the circuit depth increases from 2 (physical circuit depth) to $23(L_{y}-1)+$ 2 (compiled circuit depth) after this recompiling. The physical circuit depths and compiled circuit depths for several 2D quantum circuits studied in Fig. 3–Fig. 6 are provided in Table 1.
@@
Table 1: Physical circuit depths and compiled circuit depths for several 2D quantum circuits studied in Fig. 3–Fig. 6. Here, a 2D quantum circuit is arranged in a square lattice with the number of qubits $N=L_{x}\times L_{y}$.
@@
B.5 Two-dimensional parametrized quantum circuit (PQC-2D)
@@
Appendix C Simulation accuracy of the pTEBD algorithm for a single instance of random parameters

### v2

@@ Appendix A Details of the quantum circuits adopted in the benchmarking simulations
A.2 One-dimensional random quantum circuit (RQC-1D)
@@
where $\sigma^{x}_{r}$, $\sigma^{y}_{r}$, and $\sigma^{z}_{r}$ are Pauli matrices, and $\alpha_{r}^{(l)}$, $\theta_{r}^{(l)}$, and $\phi_{r}^{(l)}$ are uniformly distributed real random numbers. After applying $U_{r}^{(l)}$ to all qubits in the $l$th circuit layer, the quantum entanglement is generated by subsequently applying a layer of controlled-Z (CZ) gates, completing one physical layer of the circuit construction. The CZ gates apply to $(0,1),(2,3),\cdots,(N-3,N-2)$ qubit pairs for odd physical layers and on $(1,2),(3,4),\cdots,(N-2,N-1)$ qubit pairs for even physical layers. The total number $N$ of qubits is restricted to odd, and the total number $D$ of physical circuit layers is restricted to even. Note that these RQCs (including the RQC-2D introduced below) are different from those used in Google’s quantum supremacy experiment [2], in which fSim gates are adopted to generate larger entanglement. More detailed discussion of these two kinds of RQCs can be found in Ref. [7, 54].
@@
A.3 One-dimensional parametrized quantum circuit (PQC-1D)
@@
A.4 Two-dimensional random quantum circuit (RQC-2D)
RQC-2D is the direct extension of RQC-1D on a 2D square lattice with the number of qubits $N=L_{x}\times L_{y}$. The physical circuit layers are applied in an ABCDABCD $\cdots$ pattern, as illustrated in Fig. 9, where each rectangle indicates the qubits on which a CZ gate acts. Additionally, single-qubit random gates $\{U_{r}^{(l)}(\theta_{ij}^{(l)})\}$ are applied to all qubits in each physical circuit layer, similar to the case of RQC-1D. The number of physical circuit layers is restricted to a multiple of 4. However, there are no specific restrictions on $L_{x}$ and $L_{y}$. In cases when $L_{x}$ ($L_{y}$) is odd, no two-qubit gates are applied on qubits at the right and left (top and bottom) edges in physical circuit layers C and D (A and B), respectively, as shown in Fig. 9 for $L_{x}=L_{y}=5$.
@@ A.4 Two-dimensional random quantum circuit (RQC-2D)
In contrast to 1D quantum circuits, where qubits naturally align in a 1D path suitable for the MPS representation, a mapping of qubits for 2D quantum circuits to a 1D path suitable for the MPS representation must be determined. Here, we choose the path shown in Fig. 10(a) to ensure that two-qubit gates in physical circuit layers A and B act on neighboring qubits in the sense of the MPS 1D path. Under this mapping, the two-qubit gates in physical circuit layers C and D become long-distance gates, necessitating a recompilation of these two physical circuit layers for MPS-based simulations to ensure that gates apply only to neighboring qubits on the MPS 1D path. As an example, the recompiling procedure for the $L_{x}=L_{y}=4$ case is illustrated in Fig. 10(b). Note that the circuit depth increases from 2 (physical circuit depth) to $23(L_{y}-1)+$ 2 (compiled circuit depth) after this recompiling. The physical circuit depths and compiled circuit depths for several 2D quantum circuits studied in Fig. 4–Fig. 7 are provided in Table 1. Although we focus on the square lattice, an extension to other lattices such as the Bristlecone lattice used in Ref. [2] is straightforward.
@@
Table 1: Physical and compiled circuit depths for several 2D quantum circuits studied in Fig. 4–Fig. 7. Here, a 2D quantum circuit is arranged in a square lattice with the number of qubits $N=L_{x}\times L_{y}$.
@@
A.5 Two-dimensional parametrized quantum circuit (PQC-2D)
@@
Appendix B Simulation accuracy of the pTEBD algorithm for a single instance of random parameters
@@
Appendix C Additional benchmark test of pTEBD
Quantum Fourier transformation (QFT) [63] is a fundamental algorithm in fault-tolerant quantum computing (FTQC) and is typically integrated into other FTQC algorithms, such as Shor’s algorithm [64] and the quantum phase estimation algorithm [65]. In this appendix, we simulate the QFT applied to a random state using both the sequential MPS algorithm and the pTEBD algorithm to assess the accuracy and scalability of the pTEBD algorithm in the simulation of more irregularly structured quantum circuits compared with those studied in Sec. IV.2.
We initialize a random MPS with bond dimension $\chi_{0}=10$ and apply a QFT circuit to this random state. The QFT circuit is generated and further compiled to a qubits layout with linear connectivity using Qiskit [66]. For the compilation, the transpile() function is used with the single-qubit rotation gate $U(\theta,\phi,\lambda)$ and the controlled-NOT gate CX as the basis gates, and with the optimization level 3. The physical circuit depths and compiled circuit depths of the QFT circuits for different system sizes are summarized in Table 2. We then employ the sequential MPS algorithm and the pTEBD algorithm to simulate these compiled circuits.
Table 2: Physical and compiled circuit depths for QFT circuits with various numbers ($N$) of qubits.
$N$ physical circuit depth compiled circuit depth
16 136 946
20 210 1482
24 300 2219
28 406 3034
32 528 3882
48 1176 8430
64 2080 14670
80 3240 22514
96 4656 31493
First, we examine the simulation accuracy. For this purpose, we fix the number of qubits $N=16$ with the corresponding compiled circuit depth being 946. The results of the simulation accuracy, quantified by the wavefunction fidelity $\mathcal{F}$ in Eq. (38), are shown in Fig. 12. We observe that the wavefunction fidelity in the pTEBD simulation always approximates very accurately the wavefunction fidelity obtained from the sequential MPS simulation, even in the cases with relatively small bond dimensions. This indicates that the accuracy of the pTEBD algorithm is comparable to that of the sequential MPS algorithm, as in the cases for more regularly structured quantum circuits before recompilation, studied in Fig. 4.
Figure 12: Wavefunction fidelity $\mathcal{F}$ as a function of MPS bond dimension $\chi$ in simulations of the QFT circuit with $N=16$ using the sequential MPS algorithm and the pTEBD algorithm. We set $g=0$ for the pTEBD simulations. The results for each $\chi$ are obtained from the average over 10 different random initial states.
Furthermore, we compare the performance of these two algorithms in simulating QFT circuits. Figure 13 shows the elapsed time per compiled circuit layer against the number $N$ of qubits. The elapsed time per layer increases linearly with $N$ in the sequential MPS simulations, while it remains approximately constant, especially for larger $N$, in the pTEBD simulations. The very slowly increase in elapsed time per layer in the pTEBD simulations might be due to the increase in inter-node communications, as we always distribute 16 qubits to each node in these simulations. This performance analysis demonstrates that the pTEBD algorithm can also achieve good weak scaling even when simulating more general, unstructured circuits, consistent with the results for RQC-2D and PQC-2D in Figs. 6(c) and 6(d), where the quantum circuits after recompilation are not perfectly regularly structured.
Figure 13: Elapsed time per circuit layer versus the number $N$ of qubits obtained using the sequential MPS algorithm (open circles) and the pTEBD algorithm (crosses), with a fixed MPS bond dimension $\chi$ in the simulations of QFT circuits. For the pTEBD simulations, we set $g=0$ and distribute 16 qubits to each node.